# Discovery of MgS and NaS in the Interstellar Medium and tentative detection of CaO

### v1

@@ 1 Introduction
Metals111Metals: all elements heavier than hydrogen and helium. constitute a significant component of the interstellar medium (ISM) and they are known to be heavily depleted onto dust grains in dense molecular clouds (Field, 1974; Savage & Sembach, 1996; Savaglio et al., 2003; Jenkins, 2009; De Cia et al., 2016; Roman-Duval et al., 2021; Konstantopoulou et al., 2023, 2024). These dust grains are mainly composed by graphites and silicates, (Laor & Draine, 1993; Jaeger et al., 1994; Dorschner et al., 1995; Mutschke et al., 1998; Tielens, 1998; Fabian et al., 2001), with differences in the chemical composition depending on the type of astronomical source (Köhler & Li, 2010). Graphites are mainly composed of carbon (Dartois, 2019). However, silicates present heavier metal elements. We typically consider seven types of silicates: olivine (whose composition is Mg2xFe2-2xSiO4), forsterite (Mg2SiO4), fayalyte (Fe2SiO4), pyroxene222Pyroxene could also present a sodium/aluminum composition of the form Na0.5Al0.5SiO3 (Min et al., 2007). (MgxFe1-xSiO3), hematite (Fe2O3), gehlenite (Ca2Al2SiO7), and enstatite (MgSiO3).
@@ 2 Observations
We have analysed the new ultrasensitive, broadband spectral survey towards G+0.693 (Rivilla et al., 2023; Sanz-Novo et al., 2023). These observations were carried out with the Yebes 40 m (Guadalajara, Spain) IRAM 30 m (Granada, Spain) and APEX 12 m (Atacama, Chile) radiotelescopes in position-switching mode. As central coordinates, we used , , with the off position shifted by and .
@@ 3.1 Detection of sodium sulfide (NaS)
Nine rotational transitions of sodium sulfide (NaS) have been covered within our spectroscopic survey towards G+0.693. Five out of nine transitions appear unblended while the remaining four are either heavily blended with more abundant species previously identified towards G+0.693 or fall within the noise. In Figure 1 and Table 2, we present the observed spectra and the spectroscopic information of the five unblended transitions of NaS. In Table 2, we show that each observed line is produced by the blend of two different lambda states of the same transition, which are labeled by the parity notation e and f. Furthermore, all transitions detected toward G+0.693 correspond to the state. This is because the ground state of NaS is inverted, i.e. the = 3/2 component lies lower in energy than the = 1/2 sublevel (Li & Ziurys, 1997). This explains why the transitions of NaS in the = 3/2 sublevel are easier to detect than the ones in the = 1/2 sublevel. The big absorption feature seen at 79.238424-79.238730 GHz in the third panel of Figure 1, may be due to a molecular transition affected by strong non-LTE effects, or to emission in the OFF position used in our observations.
@@
We note that five unblended transitions are enough to claim the detection of metal-bearing molecules. As examples, we refer to the recent discovery of FeC and SiP in the complex and line rich spectra of the evolved star IRC+10216 (Koelemay et al., 2022; Koelemay & Ziurys, 2023), with respectively three and four rotational transitions reported. In addition, the NaS transitions shown in Figure 1 and Table 2, correspond to the +1 progression of NaS rotational lines that was covered in our observations and that appear unblended. Finally, we note that we have thoroughly analyzed the emission from all the molecules included in the CDMS (Endres et al., 2016), JPL (Pickett et al., 1998), and Lille333See https://lsd.univ-lille.fr/. spectroscopic databases in the vicinity of the target lines, and no transitions from other molecular species besides NaS are able to explain the observed spectral features.
@@
The physical parameters of the emission of NaS are reported in Table 1 and were derived using Madcuba-Autofit on the unblended transitions shown in Figure 1 that take into account the contribution of all the molecules already identified and fitted in the molecular cloud to ensure that the observed spectral features correspond to sodium sulfide. To allow the convergence of the Autofit, we fixed the FWHM value to 15 km s-1, which reproduces well the profile of the unblended lines, and left the molecular column density (), the excitation temperature () and the central radial velocity () as free parameters. Note that the molecular line emission measured towards G+0.693 presents typical linewidths of 15-20 km s-1 (see Requena-Torres et al., 2006; Zeng et al., 2018; Jiménez-Serra et al., 2020; Rivilla et al., 2020). The best LTE fit is shown in Figure 1 and provides a velocity of km s-1, which is also consistent with the values found in G+0.693 for other molecular species. The excitation temperature is K, while the derived column density is cm-2. Assuming a H2 column density of cm-2 inferred from C18O emission at the velocity component of G+0.693 (see Martín et al., 2008)444The uncertainty in the H2 column density toward G+0.693 is assumed to be of 15% its value., the NaS column density translates into a molecular abundance with respect to H2 of towards G+0.693. However, this value should be considered as a lower limit since NaS may be produced in a region smaller than that probed by C18O (see a detailed discussion in Section 4.1).
| Frequency | Transition | log I | Eu | rms | S/Na | Blending | |||
@@
| 91.425124 | 13/2 11/2 | e | -1.9179 | 17.5 | 1.7 | 46 | 5.0 | Unblended* | |
| 91.425530 | 13/2 11/2 | f | -1.9179 | 17.5 | Unblended* | ||||
| 103.610186 | 15/2 13/2, | f | -1.7576 | 22.4 | 2.2 | 45 | 3.9 | Unblended* | |
| 103.610709 | 15/2 13/2, | e | -1.7576 | 22.4 | Unblended* | ||||
| 79.238424 | 11/2 9/2, | f | -2.1044 | 13.1 | 1.9 | 43 | 3.8 | Unblended* | |
| 79.238730 | 11/2 9/2, | e | -2.1044 | 13.1 | Unblended* | ||||
| 140.153378 | 21/2 19/2, | e | -1.3831 | 40.7 | 1.4 | 30 | 5.0 | Unblended* | |
| 140.154336 | 21/2 19/2, | f | -1.3831 | 40.7 | Unblended* | ||||
| 152.329716 | 23/2 21/2, | f | -1.2835 | 48.0 | 1.0 | 23 | 5.3 | Unblended* | |
| 152.330848 | 23/2 21/2, | e | -1.2835 | 48.0 | Unblended* | ||||
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithms of the integrated intensity at 300 K (log I), energies of the upper levels (Eu), noise level (rms), integrated area (), and signal to noise ratio (S/N) of each transition. The last column indicates the line blending with other molecular species. Unblended lines refer to those transitions that are not contaminated by the emission from other species. We also use an asterisk symbol for those lines that are (auto)blended with another transition of NaS.
We have also carried out a rotational diagram analysis (see Figure 2) of the emission of NaS with Madcuba considering the observed velocity-integrated intensity over the linewidth (Rivilla et al., 2021a). As described in San Andrés et al. (2024b), MADCUBA’s rotational diagram considers the predicted emission from all of the molecular species previously detected in the source, and accounts for the blending in each transition by subtracting the predicted line profiles of the blending species from the observed spectrum. In this way, the rotational diagram can be derived using an unblended dataset of molecular transitions for the species of interest. For NaS, the detected molecular lines are basically unblended, and hence the subtraction of the blending contribution is minimal. As shown in Figure 2, the derived physical parameters of NaS are consistent with the ones obtained with the Autofit analysis: K and cm-2.
@@ 3.2 Detection of magnesium sulfide (MgS)
Following the previous methodology, we have covered up to twelve rotational transitions of magnesium sulfide (MgS) within our spectroscopic survey. Four transitions out of the 12 are unblended, while three other lines appear blended with other molecular species but reproduce almost perfectly the observed spectra (see Figure 3 and Table 3). The four remaining ones are heavily blended with other molecules towards G+0.693 or lie within the noise and are therefore not shown.
@@
| (GHz) | () | () | (K) | (mK) | (mK km s-1) | ||
| 160.10545 | 10-9 | -1.0416 | 42.1 | 1.9 | 108 | 12.1 | Unblended |
| 112.08563 | 7-6 | -1.4779 | 21.4 | 1.2 | 90 | 13.0 | s-CH2CHOH |
| 224.10317 | 14-13 | -0.6568 | 80.3 | 1.0 | 79 | 20.4 | Unblended |
| 240.09612 | 15-14 | -0.5831 | 91.7 | 0.8 | 67 | 20.2 | Unblended |
| 80.06513 | 5-4 | -1.9029 | 11.5 | 1.7 | 59 | 5.1 | Blended: t-C2H3CHO |
| 48.04067 | 3-2 | -2.5595 | 3.9 | 0.6 | 25 | 4.9 | OC33S and n-C3H7CN |
| 32.02744 | 2-1 | -3.085 | 2.3 | 0.3 | 12 | 3.5 | Unblended |
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithm of the integrated intensities at 300 K (log I), energies of the upper levels (Eu), noise level (rms), area (), and signal to noise ratio (S/N) of each transition. In the last column, we also indicate the line blending and the species with which the line is blended. “Unblended” lines refer to those transitions that are not blended with other molecular species.
@@ 3.3 Tentative detection of calcium oxide (CaO)
We have covered seven transitions of CaO in the spectroscopic survey of G+0.693. Only two out of seven transitions observed are unblended while two other are partially blended with transitions from other molecules although the total predicted line profiles reproduce nicely the observed spectra (see Figure 5 and Table 4). It is worth noting that transition 10-9 at 265.415153 GHz (fourth panel) could be blended with other unidentified species. Due to the limited number of unblended transitions, although the model of CaO can be fitted to the residual of the emission after subtracting the emission of all other modeled species, it cannot be considered conclusive and therefore the detection is claimed to be only tentative. The rest of transitions covered within the survey lie within the noise and therefore are not shown here.
@@
| (GHz) | (- ) | () | (K) | (mK) | (mK km ) | |||
| 132.737155 | 5-4 | 0 | -1.0466 | 19.0 | 0.9 | 68 | 13.9 | Blended: Cyclopropenone |
| 238.887128 | 9-8 | 0 | -0.3326 | 57.1 | 2.4 | 59 | 6.4 | Unblended |
| 106.192587 | 4-3 | 0 | -1.3290 | 12.7 | 1.7 | 54 | 5.5 | Blended: CH3OCHO |
| 265.415153 | 10-9 | 0 | -0.2129 | 69.7 | 0.7 | 47 | 17.2 | Unblended |
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithm of the integrated intensities at 300 K (log I), energies of the upper levels (Eu), noise level (rms), area (), signal to noise ratio (S/N) of each transition. In the last column, we indicate line blending and the species with which the line is blended. “Unblended” lines refer to those transitions that are not contaminated by other molecular species.

### v2

@@ 1 Introduction
Metals111We refer to metals as defined in chemistry, i.e. as every substance with high electrical conductivity, luster, and malleability, which readily loses electrons to form positive ions. These include alkali metals such as Li, Na, K, Rb, Cs and Fr, and alkaline earth metals such as Be, Mg, Ca, Sr, Ba and Ra (Atkins et al., 2018). constitute a significant component of the interstellar medium (ISM) and they are known to be heavily depleted onto dust grains in dense molecular clouds (Field, 1974; Savage & Sembach, 1996; Savaglio et al., 2003; Jenkins, 2009; De Cia et al., 2016; Roman-Duval et al., 2021; Konstantopoulou et al., 2023, 2024). These dust grains are mainly composed by graphites and silicates, (Laor & Draine, 1993; Jaeger et al., 1994; Dorschner et al., 1995; Mutschke et al., 1998; Tielens, 1998; Fabian et al., 2001), with differences in the chemical composition depending on the type of astronomical source (Köhler & Li, 2010). Graphites are mainly composed of carbon (Dartois, 2019). However, silicates present heavier metal elements. We typically consider seven types of silicates: olivine (whose composition is Mg2xFe2-2xSiO4), forsterite (Mg2SiO4), fayalyte (Fe2SiO4), pyroxene222Pyroxene could also present a sodium/aluminum composition of the form Na0.5Al0.5SiO3 (Min et al., 2007). (MgxFe1-xSiO3), hematite (Fe2O3), gehlenite (Ca2Al2SiO7), and enstatite (MgSiO3).
@@ 2 Observations
We have analysed the new ultrasensitive, broadband spectral survey towards G+0.693 (Rivilla et al., 2023; Sanz-Novo et al., 2023). These observations were carried out with the Yebes 40 m (Guadalajara, Spain) IRAM 30 m (Granada, Spain) and APEX 12 m (Atacama, Chile) radiotelescopes in position-switching mode. As central coordinates, we used , , with the off position shifted by and . The location of G+0.693 corresponds to source Sgr B2M (20, 100) in Table 1 of Martín et al. (2008).
@@ 3.1 Detection of sodium sulfide (NaS)
Nine rotational transitions of sodium sulfide (NaS) have been covered within our spectroscopic survey towards G+0.693. Five out of nine transitions appear unblended from known molecular species, while the remaining four are either heavily blended with more abundant species previously identified towards G+0.693 or fall within the noise. For this analysis, we have thoroughly searched all the molecules included in the CDMS (Endres et al., 2016), JPL (Pickett et al., 1998), and Lille333See https://lsd.univ-lille.fr/. spectroscopic databases in the vicinity of the NaS target lines, and no transitions from other molecular species besides NaS are able to explain the observed spectral features. Despite this, the transitions 15/213/2 and 11/29/2 appear blended with unidentified line emission (U-line). However, as explained in Section 4.1, the level of blending is 25% and therefore, we consider them as “Unblended.
@@
In Figure 1 and Table 2, we present the observed spectra and the spectroscopic information of the three unblended transitions of NaS, together with the two lines blended with U-line emission. In Table 2, we show that each observed line is produced by two different lambda states of the same transition, which are labeled by the parity notation “e and “f. Furthermore, all transitions detected toward G+0.693 correspond to the state. This is because the ground state of NaS is inverted, i.e. the = 3/2 component lies lower in energy than the = 1/2 sublevel (Li & Ziurys, 1997). This explains why the transitions of NaS in the = 3/2 sublevel are easier to detect than the ones in the = 1/2 sublevel. The big absorption feature seen at 79.238424-79.238730 GHz in the third panel of Figure 1, may be due to a molecular transition affected by strong non-LTE effects, or to emission in the OFF position used in our observations.
@@
We note that three unblended transitions are enough to claim the detection of metal-bearing molecules (see the validity of the NaS detection in Section 4.1). As other examples, we refer to the recent discovery of FeC and SiP in the complex and line rich spectra of the evolved star IRC+10216 (Koelemay et al., 2022; Koelemay & Ziurys, 2023), with respectively three and four rotational transitions reported. In addition, the NaS transitions shown in Figure 1 and Table 2, correspond to the +1 progression of NaS rotational lines that was covered in our broadband observations across (non-continuous) 124 GHz, and that appear unblended (or blended with U-emission) according to our criteria (see Section 4.1). Because of all this, the data presented here give very strong evidence in support of the presence of NaS in the ISM.
@@
The physical parameters of the emission of NaS are reported in Table 1 and were derived using Madcuba-Autofit on the unblended transitions shown in Figure 1 that take into account the contribution of all the molecules already identified and fitted in the molecular cloud to ensure that the observed spectral features correspond to sodium sulfide. To allow the convergence of the Autofit, we fixed the FWHM value to 15 km s-1, which reproduces well the profile of the unblended lines, and left the molecular column density (), the excitation temperature () and the central radial velocity () as free parameters. Note that the molecular line emission measured towards G+0.693 presents typical linewidths of 15-20 km s-1 (see Requena-Torres et al., 2006; Zeng et al., 2018; Jiménez-Serra et al., 2020; Rivilla et al., 2020). The best LTE fit is shown in Figure 1 and provides a velocity of km s-1, which is also consistent with the values found in G+0.693 for other molecular species. The excitation temperature is K, while the derived column density is cm-2. Assuming a H2 column density of cm-2 inferred from C18O emission at the velocity component of G+0.693 (see Martín et al., 2008)444The uncertainty in the H2 column density toward G+0.693 is assumed to be of 15% its value. However note that this estimate may suffer from an additional bias by a factor of 2 because Martín et al. (2008) used the relationship C18O/ cm-2 (Frerking et al., 1982) derived from the molecular clouds in the Galactic disk, and there is a factor of 2 difference in the isotopic ratio 16O/18O measured in the Galactic Center compared to the Solar neighbourhood. This means that the abundance of NaS could be a factor of 2 higher., the NaS column density translates into a molecular abundance with respect to H2 of towards G+0.693. However, this value should be considered as a lower limit since NaS may be produced in a region smaller than that probed by C18O (see a detailed discussion in Section 4.2).
Frequency | Transition | log I | Eu | rms | d | S/Na | Blending | ||
@@
91.425124 (80)b | 13/2 11/2 | e | -1.9179 | 17.5 | 1.7 | 92 (9)c | 10.0 | Unblended* | |
91.425530 (80)b | 13/2 11/2 | f | -1.9179 | 17.5 | Unblended* | ||||
103.610186 (86)b | 15/2 13/2, | f | -1.7576 | 22.4 | 2.2 | 90 (12)c | 7.8 | Blended with U-lines* | |
103.610709 (86)b | 15/2 13/2, | e | -1.7576 | 22.4 | Blended with U-lines* | ||||
79.238424 (72)b | 11/2 9/2, | f | -2.1044 | 13.1 | 1.9 | 86 (11)c | 7.6 | Blended with U-lines* | |
79.238730 (72)b | 11/2 9/2, | e | -2.1044 | 13.1 | Blended with U-lines* | ||||
140.153378 (97)b | 21/2 19/2, | e | -1.3831 | 40.7 | 1.4 | 60 (6)c | 10.0 | Unblended* | |
140.154336 (97)b | 21/2 19/2, | f | -1.3831 | 40.7 | Unblended* | ||||
152.329716 (97)b | 23/2 21/2, | f | -1.2835 | 48.0 | 1.0 | 46 (4)c | 10.6 | Unblended* | |
152.330848 (97)b | 23/2 21/2, | e | -1.2835 | 48.0 | Unblended* |
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithms of the integrated intensity at 300 K (log I), energies of the upper levels (Eu), noise level (rms), integrated area (), and signal to noise ratio (S/N) of each transition. The last column indicates the line blending with other molecular species. “Unblended” lines refer to those transitions that are not contaminated by the emission from other known species. We also use an asterisk symbol for those lines that are (auto)blended with another transition of NaS.
We have also carried out a rotational diagram analysis (see Figure 2) of the emission of NaS with Madcuba considering the observed velocity-integrated intensity over the linewidth (Rivilla et al., 2021a). As described in San Andrés et al. (2024b), MADCUBA’s rotational diagram considers the predicted emission from all of the molecular species previously detected in the source, and accounts for the blending in each transition by subtracting the predicted line profiles of the blending species from the observed spectrum. In this way, the rotational diagram can be derived using an “unblended dataset of molecular transitions for the species of interest. For NaS, the detected molecular lines are basically unblended, and hence the subtraction of the blending contribution is minimal. As shown in Figure 2, the derived physical parameters of NaS are consistent with the ones obtained with the Autofit analysis: K and cm-2.
@@ 3.2 Detection of magnesium sulfide (MgS)
Following the previous methodology, we have covered up to twelve rotational transitions of magnesium sulfide (MgS) within our spectroscopic survey. Three transitions out of the 12 are unblended, while one is blended with a U-line and three other appear blended with known molecular species but reproduce almost perfectly the observed spectra (see Figure 3 and Table 3). The five remaining transitions are heavily blended with other molecules towards G+0.693 or lie within the noise and are therefore not shown.
@@
(GHz) | (-) | () | (K) | (mK) | (mK km s-1) | ||
160.105451 (3)b | 10-9 | -1.0416 | 42.1 | 1.9 | 108 (9)c | 12.1 | Unblended |
112.085630 (2)b | 7-6 | -1.4779 | 21.4 | 1.2 | 90 (7)c | 13.0 | s-CH2CHOH |
224.103178 (15)b | 14-13 | -0.6568 | 80.3 | 1.0 | 79 (4)c | 20.4 | Unblended |
240.096126 (3)b | 15-14 | -0.5831 | 91.7 | 0.8 | 67 (3)c | 20.2 | Unblended |
80.065135 (2)b | 5-4 | -1.9029 | 11.5 | 1.7 | 59 (12)c | 5.1 | t-C2H3CHO |
48.040669 (1)b | 3-2 | -2.5595 | 3.9 | 0.6 | 25 (5)c | 4.9 | OC33S and n-C3H7CN |
32.027444 (0.7)b | 2-1 | -3.085 | 2.3 | 0.3 | 12 (3)c | 3.5 | Blended with U-line |
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithm of the integrated intensities at 300 K (log I), energies of the upper levels (Eu), noise level (rms), area (), and signal to noise ratio (S/N) of each transition. In the last column, we also indicate the line blending and the species with which the line is blended. “Unblended” lines refer to those transitions that are not blended with other known molecular species.
@@ 3.3 Tentative detection of calcium oxide (CaO)
We have covered seven transitions of CaO in the spectroscopic survey of G+0.693. Only two out of seven transitions observed are unblended while two other are partially blended with transitions from other molecules although the total predicted line profiles reproduce nicely the observed spectra (see Figure 5 and Table 4). It is worth noting that transition 10-9 at 265.415153 GHz (fourth panel) could be blended with other unidentified species (U-species). Due to the limited number of unblended transitions, although the model of CaO can be fitted to the residual of the emission after subtracting the emission of all other modeled species, it cannot be considered conclusive and therefore the detection is claimed to be only tentative. The rest of transitions covered within the survey lie within the noise and therefore are not shown here.
@@
(GHz) | ( ) | () | (K) | (mK) | (mK km ) | |||
132.737155 (30)b | 5-4 | 0 | -1.0466 | 19.0 | 0.9 | 68 (5)d | 13.9 | Cyclopropenone |
238.887128 (46)b | 9-8 | 0 | -0.3326 | 57.1 | 2.4 | 59 (9)d | 6.4 | Unblended |
106.192587 (30)b | 4-3 | 0 | -1.3290 | 12.7 | 1.7 | 54 (10)d | 5.5 | CH3OCHO |
265.415153 (74)b | 10-9 | 0 | -0.2129 | 69.7 | 0.7 | 47 (3)d | 17.2 | Blended with U-line |
Note. — We provide the transition frequencies, quantum numbers, base 10 logarithm of the integrated intensities at 300 K (log I), energies of the upper levels (Eu), noise level (rms), area (), signal to noise ratio (S/N) of each transition. In the last column, we indicate line blending and the species with which the line is blended. “Unblended” lines refer to those transitions that are not contaminated by other known molecular species.
@@ 4 Discussion
4.1 Validity of the detections
To check the validity of the NaS and MgS detections, we have followed the criteria of Snyder et al. (2005) and Halfen et al. (2006) for the definition of an unblended line. The criteria used are the following:
2. The VLSR of the emission of NaS and MgS should lie at 69 km s-1 with a tolerance of 5 km s-1. The measured VLSR for NaS is 66.5 km s-1, while it is 69 km s-1 for MgS (see Table 1), well within those limits. Note that this VLSR tolerance is significantly narrower than the one assumed by Snyder et al. (2005) for Sgr B2(N).
3. The linewidths of the NaS and MgS lines should lie between 15 and 22 km s-1, as observed for other molecular species toward G+0.693 (see e.g. Rivilla et al., 2020; Massalkhi et al., 2023). The linewidths that best reproduce the NaS and MgS spectra are 15 km s-1 (Table 1), consistent with these values.
4. The unblended lines should not show too great intensity. As shown by the LTE fits of Figures 1 and 3 obtained with SLIM-MADCUBA, the observed spectral features are perfectly reproduced by the expected LTE intensity.
5. There must be spectral features at all favorable, physically connected transitions over a sufficiently large wavelength range. As mentioned in Sections 3.1 and 3.2, we have detected the full progression of +1 transitions of NaS and MgS covered in our broadband spectral survey across (non-continuous) 124 GHz.
6. Snyder et al. (2005) considered that if the observed feature appeared to be a blend of transitions from several species, this line was not used in the analysis. As explained in Section 3.1, the observed lines of NaS and MgS are unblended from known molecular species, although some transitions may appear blended with emission from U-species. To determine the level of contribution of the U-species, we have obtained the residuals after subtracting the LTE fit of the NaS and MgS lines, and measured the area of the residuals for the velocity range between FWHM/2, with FWHM the measured linewidth of the NaS and MgS lines (in this case, 15 km s-1 for both species). If this contribution is 25%, the transition is labeled as “Unblended in Tables 2 and 3, and therefore, they are used in our analysis and to claim a detection of these molecular species. If instead, the contribution is 25%, the transitions are labeled as “Blended with U-line in Tables 2 and 3, and they are not used to claim the detection. This is the case of e.g. the 21 line of MgS or of the 15/213/2 and 11/29/2 lines of NaS.
Following Halfen et al. (2006), we have estimated the confidence level of the NaS and MgS detections considering the “Unblended transitions. Assuming the confusion limit (see Halfen et al., 2006), the chance of finding a line at the VLSR of our source with a tolerance of 3 km s-1 (i.e. the difference in VLSR observed for NaS and MgS) is =32% assuming simple Gaussian profiles and a velocity coverage of the observed feature above the rms noise level equivalent to the FWHM at 1/3 of the peak intensity. The detection of each successive line after the initial one will have a probability (1st)(2nd)=. Consequently, three unblended lines provides a probability of chance alignment of =3.3%, which implies a confidence level of 96.7% for the detection of NaS and MgS. If four or five lines were considered (i.e. if we included the transitions blended with U-lines), we would obtain confidence levels of 99% and 99.7% for the detection of NaS and MgS, respectively. Note that in any case these are conservative estimates, since we have assumed the confusion limit in our calculations and this limit has not been reached yet in our survey toward G+0.693.